# Protocol as Poetry: Case Study on Pak’s Protocol Arts

### v1

@@ 3. Method
We employed a qualitative, grounded-theory strategy to map the complete corpus of Pak’s protocol artworks and theorise the behaviours they generate. Data collection proceeded four complementary strata: (1) On-chain evidence: smart-contract source code and transaction logs for every work, extracted via Dune Analytics and OpenSea. (2) Project interfaces: official webpages, and marketplace listings (e.g. Nifty Gateway launch page). (3) Discourse layer: journalistic reviews, critical essays, and high-signal social-media threads that register community reception. (4) Authorial voice: archived interviews and statements from Pak’s Twitter feed, providing insight into artistic intent.
Building on the manually curated corpus described above, we broadened coverage by deploying a large-language-model (LLM) agent that exhaustively scraped and consolidated public references to nine landmark protocol artworks by Pak. Guided by theoretical sampling, we subjected this enlarged dataset to iterative open coding; through constant comparison, the emerging codes crystallised into seven analytic lenses rooted in protocol-art scholarship: (1) Concept, (2) System–mechanism dynamics, (3) Participatory interaction, (4) Artist’s control, (5) Collective emergent behaviour, (6) Poetic meaning-making, and (7) Key statistics. The LLM next pre-classified each artefact under these lenses (see Appendix), after which two researchers manually verified, corrected, and enriched every entry.
@@
During focused coding, we refined properties and dimensions within each lens and annotated every project in a comparative matrix (see Table 1). Axial coding then traced connections across lenses, revealing higher-order patterns, category saturation—no new properties emerging from fresh data—signalled theoretical sufficiency, yielding the grounded model of Pak’s practice that structures the subsequent discussion.
4. Result: Characterizing Pak’s Protocol Arts
Our grounded-theory analysis reveals that Pak’s oeuvre operates as a coherent yet evolving family of protocol artworks whose aesthetic force resides less in visual form than in the programmable conditions that orchestrate collective action. Across nine landmark projects—from The Fungible and its cube-fusion economy to the auto-aggregating masses of Merge, the speech-filtering ritual of Censored, and the self-reflexive void of #404—we observe a stable configuration of traits following. Taken together, these features constitute what we term Pak’s protocol aesthetic: a design grammar that transforms blockchain infrastructure into a stage for large-scale, participatory meaning-making.
@@ 4.1. System-centric rather than object-centric composition
Pak’s practice exemplifies a shift from discrete art objects to dynamic systems. In a “systems aesthetic,” as Jack Burnham theorized, the artist designs goals, rules, and interactions rather than crafting a fixed artifact. Unlike a static object with defined boundaries, a system can evolve over time and respond to external inputs. Pak explicitly embraces this ethos: he argues that an NFT’s image is merely one element of a token’s metadata, and focusing only on the visual is “a naive frame of view”. The true medium for Pak is the underlying protocol – “the structure of the NFT market itself” – which he leverages as a creative material.
This system-centric approach is evident in works like The Fungible (2021). Rather than a singular digital image, The Fungible was presented as a three-day performative market game on Nifty Gateway. Buyers effectively co-created the outcome: they could purchase unlimited open-edition “Cube” tokens, and “the more they bought, the more unique their NFT would be”. The artwork unfolded through rules of supply and demand, competition, and reward, blurring the line between artistic composition and economic system. Similarly, Pak’s earlier collection X (2020) consisted of “infinite editions” available for only one day, such that scarcity was determined by time and participation rather than a predefined edition size. In all these cases, what constitutes the artwork is not a static file but the entire process and protocol: the contracts, algorithms, and user interactions that generate the final outputs. This aligns with Burnham’s notion that in system-based art “all phases of the life cycle of a system are relevant” and there is “no end product that is primarily visual”. Pak’s art finds its form in the living system of relationships and behaviors, not in a singular immutable image.
@@ 4.2. Autonomous governance for open-ended control
Many of Pak’s projects foreground autonomous processes and decentralized control, resonating with concepts of generative autonomy in both blockchain and artificial life. On the blockchain, smart contracts act as self-governing “digital physics” – immutable rule sets that execute without centralized oversight. This allows artworks to behave like autonomous entities. For example, Merge (2021) implemented a self-contained rule whereby all purchased units automatically “merged” into singular tokens in each wallet at the close of the sale, without any manual intervention. The final form emerged from the contract’s code and collective user actions, illustrating open-ended control dictated by an algorithm. Pak often cedes such control to the system itself or the community, allowing outcomes beyond a single author’s full prediction.
This ethos connects to emerging “autonomous worlds” in crypto art, and is exemplified by projects like Primavera De Filippi’s Plantoid, an art piece that lives on a DAO and “reproduces itself” via smart contract governance. In Plantoid, people who hold “seed” NFTs fund and vote on the creation of new instances, effectively making the artwork a self-propagating entity. Pak’s work incorporates similar principles of shared or automated governance. In Censored (2022), the one-of-one Clock NFT was auctioned not to an individual but to AssangeDAO, a collective of over 10,000 people pooling funds to bid. A decentralized community thus became the owner, blurring the line between patron and curator through collective governance. Moreover, Pak’s burn.art platform runs as an ongoing autonomous mechanism: anyone can send an NFT to be burned at any time in exchange for Ash, which in turn can be used in future projects. This “never-ending ‘Buy–Burn’ game” can continue indefinitely by design, illustrating how the artwork’s evolution is handed over to participant actions and blockchain logic. Such autonomous, perpetual systems align with the idea of art as an “unstoppable” process — a “complex living system” on chain that “no single entity can halt”. Pak’s protocol art thus moves toward open-endedness, where creation and control are distributed across code and community rather than centralized in the artist.
@@ 4.3. Distributed agency and communal authorship
Similarly, Censored (2022) turned its audience into co-authors. During the 48-hour open edition, nearly 14,000 messages were tokenized by users and added to the Censored collection. Each participant essentially created a piece of the overall artwork by embedding a personal message on the blockchain. The final collection is thus a mosaic of voices from the community (“Pak & Assange & You,” as the creators billed it), rather than a singular narrative. Even Merge can be seen as distributing authorship: almost 30,000 buyers collectively determined the supply and composition of that work’s outcome, making the “most expensive NFT” a product of crowd dynamics as much as Pak’s initial concept. This radical dispersal of agency recalls Umberto Eco’s idea of the “open work,” extended into the tokenized realm – the artwork is an evolving network of contributions. By building frameworks that require audience input to complete the work, Pak positions the community as co-creators. The artistic authorship becomes plural and emergent from the group, rather than originating solely from the individual artist.
@@ 4.4. Temporal dynamism and lifecycle aesthetics
Pak’s works are not static snapshots; they unfold over time, embracing change, ephemerality, and defined life cycles as core aesthetic elements. In classical systems art, the “consistency of a system may be altered in time and space” and all stages of its life cycle become material for the artwork. Pak’s projects exemplify this temporal dynamism. Lost Poets (2021) was structured as a narrative in multiple acts – from the initial sale of “Pages,” to a delayed Reveal where pages transformed into Poet NFTs, through daily drops of special “Origin” Poets, and onward towards a planned “Twist” and final Epilogue after 365 days. The piece lives and evolves across an entire year, its full meaning only emerging gradually as collectors engage over time. This built-in lifecycle – birth (minting), maturation (naming and interacting), and an end-state after one year (when burning for Ash is enabled) – turns time into an artistic medium.
@@
Other Pak works similarly rely on temporal structure. Censored’s Clock is explicitly a time-based artwork: a counter of Assange’s days in prison that updates daily, tethering the piece’s content to the passage of time in the real world. The open edition portion of Censored was also time-limited to 48 hours, highlighting the urgency and performative moment of the audience’s participation. In X (2020), each NFT “moment” was only available during a one-day window, making “one single day of infinite existence” the defining limit of each edition. Scarcity and content in X were thus time-contingent: the fewer people who acted within the day, the fewer copies existed. Even Pak’s market dynamics often have temporal phases (for instance, The Fungible’s escalating price tiers each day). By designing pieces with evolving states and deadlines, Pak “recreates the cycle of life” in digital form, acknowledging that in a systemic artwork, impermanence and transformation are not byproducts but the very substance of the art.
4.5. Economic-driven engagement

### v2

@@ 3. Method
To investigate the emerging characteristics of protocol art, we conducted an annotated portfolio analysis (Gaver and Bowers, 2012) of Pak’s nine selected landmark protocol artworks, listed in Table 1. We compiled data from four complementary strata: (1) On-chain evidence—we examined smart-contract source code and transaction logs for every major Pak project (using resources like Etherscan [^9] https://etherscan.io/ and Dune Analytics [^10] https://dune.com/); (2) Project interfaces—we reviewed official launch webpages, marketplace listings (e.g., Nifty Gateway [^11] https://niftygateway.com/ drop pages), and any interactive frontend elements; (3) Discourse layer—we gathered journalistic articles, critical essays, and high-signal social media threads (e.g., Twitter analyses) that captured community reception and interpretation; (4) Authorial voice—we collected Pak’s own statements from archived interviews and tweets to understand the artist’s intent in each project.
Corpus Selection.
We assembled a portfolio of Pak’s protocol artworks using purposive sampling with explicit inclusion criteria: (1) the work’s primary medium is smart contract mechanics, not merely NFT-based distribution of pre-existing digital images; (2) the work involved documented collective participation, operationalized as more than 100 unique wallet addresses interacting with the contract; (3) the work generated substantial critical discourse, operationalized as at least three independent published analyses in art journalism, academic venues, or high-engagement social media threads. Nine works met all criteria (Table 1). We excluded Archillect (Pak’s 2014 AI curation bot, which predates blockchain and lacks protocol art characteristics) and minor releases without significant documentation or collective participation.
LLM-Assisted Data Gathering.
We broadened coverage by deploying a large-language-model agent (ChatGPT-o3, March 2024 version) to aggregate publicly available references to each artwork. The prompt template was: “Compile all publicly documented information about [artwork name] by Pak, including: official announcements, smart contract mechanics, critical reception, and participation statistics. Cite sources.” LLM outputs served as initial summaries requiring human verification. The first author independently verified all factual claims against primary sources: smart contract code on Etherscan, official project websites, archived tweets, and transaction records on Dune Analytics. Discrepancies were resolved in favor of primary sources.
Thematic Analysis.
Guided by theoretical sampling, we subjected the aggregated dataset to iterative open coding. Initial codes were generated inductively from the first three works (X, The Title, The Fungible), yielding 23 preliminary codes (e.g., “time-based scarcity,” “collector-as-author,” “economic gamification”). These codes were refined through constant comparison with subsequent works, guided by prior scholarship on conceptual art (LeWitt, 1967), systems art (Burnham, 1968), generative art (Boden and Edmonds, 2009), participatory art (Bishop, 2023), and interactive art (Seevinck, 2017). After analyzing Merge (the fourth work), codes stabilized into seven analytical categories: (1) Concept; (2) System–mechanism dynamics; (3) Participatory interaction; (4) Artist’s control; (5) Collective emergent behavior; (6) Poetic meaning-making; (7) Key statistics; and (8) Annotation. The LLM agent assisted in initial classification under the first seven lenses (see Appendix A for compiled entries), after which one human researcher manually verified, corrected, and enriched every entry with additional details from primary sources. The eighth lens, Annotation, was composed entirely by the human researcher, mapping each work to the seven characteristics identified in Section 4 and providing interpretive commentary grounded in the verified data.
@@
During focused coding, we refined the properties and dimensions within each lens and constructed a comparative matrix of all projects (distilled in Table 1). We then performed axial coding to trace connections across lenses and identify higher-order themes and consistent patterns. This process yielded a grounded characteristic model of Pak’s protocol art practice that structures the discussion that follows.
Project (Year) Description Participatory Interaction System Mechanism Artist’s Control Participants / Scale Collective Emergent Behavior
X (2020) Introduced time-based scarcity with a 24-hour unlimited open edition, shifting value from object to moment. 24-hour open-edition mint; collectors could mint unlimited copies within the time window. Time-based scarcity: supply fixed by duration, not by a preset count. None after launch (rules immutable once deployed). 15 NFT pieces; 13 open editions (61 editions minted) + 2 one-of-one auctions (unique pieces). Hundreds of wallets participated. “FOMO”-driven rush as deadline approached; final supply decided by community minting behavior.
The Title (2021) Identical image sold under differently priced on-chain names, showing metadata (title/pricing) can dictate value more than pixels. Fixed-price and auction sales of visually identical NFTs, each bearing a different on-chain title and price; one piece intentionally left unsold. Value experiment: image constant, metadata (title + scarcity) varies per token. Only predetermined prices/titles; tokens immutable after mint. 9 token types (6 sold via various formats, 1 remained unsold, plus gifts); ~300 total editions; dozens of collectors across formats. Community debate on what confers an NFT’s value (image vs. narrative vs. price); social media buzz about who obtained “The Expensive,” etc.
The Fungible (2021) Sotheby’s open-edition turned fungibility into spectacle: thousands of interchangeable cubes questioned uniqueness. Hybrid drop: open-edition “Cube” tokens minted in large quantities + puzzle/leaderboard rewards during a Sotheby’s event. Fungible cubes could be aggregated into higher-tier NFTs; gamified rewards for top accumulators and puzzle solvers. Rules and reward tiers fixed in contract; artist manually validated puzzle answers and distributed bonus NFTs. 23,598 cubes bought by 3,080 wallets $\rightarrow$ yielded 6,165 unique NFTs (after merging by quantity). Competitive accumulation (to get larger cube clusters); cooperative puzzle-solving; viral social media tasks as part of the game.
burn.art (2021) / \$ASH Perpetual “creation-through-destruction”: burn any NFT to mint the social currency \$ASH, gamifying scarcity. Ongoing: anyone can burn any NFT on the platform and receive \$ASH (ERC-20 token) via a bonding-curve formula. “Creation via destruction”: burn an NFT, get fungible \$ASH; token yield halves at supply milestones. Contract rules fixed; artist later released drops purchasable only with \$ASH, influencing demand. Thousands of NFTs burned; thousands of \$ASH holders emerged. Emergent burn economy: holders strategized which NFTs to sacrifice; a subculture of “burners” formed, embracing destruction as creation.
Lost Poets (2021–22) Multi-stage strategy game of 65,536 AI-generated Pages evolving or burning into Poets; collectors co-wrote an ever-shifting narrative. Multi-stage narrative game: 65,536 “Page” NFTs sold; holders could burn Pages for \$ASH or transform Pages into AI-generated “Poet” NFTs; on-chain naming of Poets. Multi-act smart contract: burn-vs-mint branching, periodic airdrops of 1,024 special “Origin” Poets to top participants; progressive trait reveals over time. Pak triggered each Act (I–IV) via contract calls and airdrops, but could not reverse user choices or outcomes once each phase started. Approx. 8,000 unique holders; resulted in approx. 28,000 Poet NFTs (others remained as Pages or were burned). Rich strategy debates (whether to keep Pages, burn for \$ASH, or transform to Poets); “naming guilds” formed to coordinate thematic naming; puzzle-hunting for hidden clues in traits.
Invisible Mechanism (a.k.a. Hate) (2021) Premiered a move() contract granting unilateral token relocation by the artist—an audacious take on creator sovereignty. Airdrop performance: 30 of Pak’s outspoken critics were invited to provide wallet addresses and received non-transferable “Hate” NFTs. Custom Move contract: tokens soul-bound to recipient (non-transferable), with an admin-only move() function letting the artist seize or relocate tokens at will. Absolute control: artist could reclaim or change tokens at any time; recipients had no control. 30 forced participants (recipients of the airdrop). Sparked wide discussion on ownership and power asymmetry in NFTs; recipients tried and failed to remove tokens; community debated the ethics of an all-powerful contract.
Merge (2021) 312,686 “mass units” sold that algorithmically fused into larger single tokens after transfers—ownership as live sculpture. 48-hour open sale of “mass units”; every buyer received a single NFT whose “mass” automatically increased as they bought or acquired more units. Self-merging deflationary tokens: all units in one wallet fused into one token; built-in leaderboard tiers (numbered 1–4 by color, with a dynamic “Alpha” class for the top holder). Rules fixed in code; artist only involved post-sale via orchestrating an ARG and acknowledging top holders. 28,983 buyers; 312,686 total mass units sold $\rightarrow$ resulted in 28,983 NFTs at sale end (one per buyer). Frenzied “mass hoarding” competition; after sale, mergers between wallets reduced token count (“extinction game”); a DAO (“Internship”) even formed to collectively purchase the largest mass (Alpha).
Censored (2022) With Julian Assange: DAO-funded 1/1 “Clock” counting prison days plus open-edition Messages—provenance weaponized for protest. Two-part project: (a) a one-of-one Clock auction via DAO, (b) a 48-hour open-edition Message mint where anyone could submit any text at any price. (a) Clock is a dynamic counter tied to Assange’s status (days in prison, now days free since his June 2024 release); (b) Message NFTs initially showed each text as blacked-out and were revealed upon Assange’s release. Clock runs autonomously; reveal was bound to Assange’s release (triggered June 2024); messages are immutable once written (Pak cannot alter or censor them). Clock was won by AssangeDAO (10,000+ members collectively own it); 29,766 Message NFTs were minted (one per wallet). Grass-roots mobilization: a global DAO formed to bid for Clock; thousands minted censored messages creating a “censored chorus.” Shared anticipation of the potential reveal unites participants; project bridged crypto art and political activism.
Not Found (#404) (2023) A 1-of-1 NFT deliberately devoid of metadata (token ID 404), transforming “page not found” into a meditation on digital absence. Single 1-of-1 charity auction; NFT token has no image or metadata (only token ID “404” visible). “Deliberate absence”: token’s metadata contract returns the string “In memory of the absent” instead of standard metadata, referencing the HTTP 404 error via its token ID. Designed absence: artist coded the contract to return only a memorial string in place of metadata; the metadata contract is technically mutable but left empty by intent. 1 owner (noted collector WhaleShark won the auction). Community reflection on owning “nothing”; sparked discourse on whether a void can be art. Served as a memorial tribute (for the late Alotta Money) that many in the community shared and discussed.
Table 1. Comparative overview of selected Pak projects
4. Results: Characterizing Pak’s Protocol Arts
Our analysis reveals that Pak’s work forms a coherent yet evolving family of protocol artworks. Their aesthetic power lies not in visual form but in the programmable conditions that orchestrate collective action. Across nine landmark projects—X (2020) [^12] https://opensea.io/collection/x-by-pak, The Title (2021) [^13] https://niftygateway.com/collections/thetitle, The Fungible (2021) [^14] https://www.sothebys.com/en/digital-catalogues/the-fungible-collection-by-pak, burn.art (2021) [^15] https://burn.art, Lost Poets (2021) [^16] https://lostpoets.xyz/, Invisible Mechanism (2021), Merge (2021) [^17] https://www.niftygateway.com/collections/pakmerge/, Censored (2022) [^18] https://opensea.io/collection/censored-pak-assange, and Not Found (2023) [^19] https://opensea.io/collection/in-memory-of-the-absent, we observe a noticeable set of characteristics. Together, these characteristics constitute what we call Pak’s protocol aesthetic—a design grammar that transforms blockchain infrastructure into a stage for large-scale participatory meaning-making. We articulate this aesthetic through seven interlinked characteristics, elaborated in the subsections below.
@@ 4.1. System-centric rather than object-centric composition
Pak’s practice exemplifies a shift from creating discrete art objects to designing dynamic systems. In a “systems aesthetic” as theorized by Jack Burnham in the late 1960s (Burnham, 1968), the artist defines goals, rules, and interactions rather than crafting a fixed artifact. Unlike a static painting or sculpture with defined forms, a system-based artwork can evolve over time and respond to external inputs. Pak explicitly embraces this ethos. Pak has argued in interviews and on social media that an NFT’s image is merely one element of a token’s metadata, and that focusing only on the visual aspect is “a naïve frame of view” (implying that the structure and behavior encoded in the token are equally part of the art). The true medium for Pak is the underlying protocol—“the structure of the NFT market itself(Marraccini, 2021), which Pak leverages as a creative material (Poposki, 2024b).
This system-centric approach is evident in works like The Fungible (2021). Rather than a singular digital image, The Fungible was presented as a three-day performative market game on Nifty Gateway. Buyers effectively co-created the outcome: they could purchase unlimited open-edition “Cube” tokens, and “the more they bought, the more unique their NFT would be”. The artwork unfolded through rules of supply and demand, competition, and reward, blurring the line between artistic composition and economic system. Similarly, Pak’s earlier collection X (2020) consisted of “infinite editions” available for only one day, such that scarcity was determined by time and participation rather than a predefined edition size. In all these cases, what constitutes the artwork is not a static file but the entire process governed by protocol: the contracts, algorithms, and user interactions that generate the final outputs. This aligns with Burnham’s notion that in system-based art “all phases of the life cycle of a system are relevant” and there is “no end product that is primarily visual” (Burnham, 1968). Pak’s art finds its form in the living system of relationships and behaviors, not in a singular immutable image.
@@ 4.2. Autonomous governance for open-ended control
Many of Pak’s projects foreground autonomous processes and decentralized control, resonating with concepts of generative autonomy in both blockchain and artificial life. On the blockchain, smart contracts act as self-governing “digital physics”—immutable rule sets that execute without centralized oversight (ludens, 2022). This allows artworks to behave like autonomous entities. For example, Merge (2021) implemented a self-contained rule whereby all purchased units automatically “merged” into singular tokens in each wallet at the close of the sale, without any manual intervention. The final form emerged from the contract’s code and collective user actions, illustrating open-ended control dictated by an algorithm. Pak often cedes such control to the system itself or the community, allowing outcomes beyond a single author’s full prediction.
This ethos connects to emerging “autonomous worlds” (ludens, 2022) in crypto art, and is exemplified by projects like Primavera De Filippi’s Plantoid, an art piece that lives on a DAO and “reproduces itself” via smart contract governance. In Plantoid, people who hold “seed” NFTs fund and vote on the creation of new instances, effectively making the artwork a self-propagating entity. Pak’s work incorporates similar principles of shared or automated governance. In Censored (2022), the one-of-one Clock NFT was auctioned not to an individual but to AssangeDAO, a collective of over 10,000 people pooling funds to bid. A decentralized community thus became the owner, blurring the line between patron and curator through collective governance. Moreover, Pak’s burn.art platform runs as an ongoing autonomous mechanism: anyone can send an NFT to be burned at any time in exchange for \$ASH, which in turn can be used in future projects. This “never-ending ‘Buy–Burn’ game” can continue indefinitely by design, illustrating how the artwork’s evolution is handed over to participant actions and blockchain logic. Such autonomous, perpetual systems align with the idea of art as an “unstoppable” process—a “complex living system” on chain that “no single entity can halt” (Hu and Fangting, 2024). Pak’s protocol art thus moves toward open-endedness, where creation and control are distributed across code and community rather than centralized in the artist.
@@ 4.3. Distributed agency and communal authorship
Similarly, Censored (2022) turned its audience into co-authors. During the 48-hour open edition, nearly 30,000 messages were tokenized by users and added to the Censored collection. Each participant essentially created a piece of the overall artwork by embedding a personal message on the blockchain. The final collection is thus a mosaic of voices from the community (“Pak & Assange & You,” as the creators billed it), rather than a singular narrative. Even Merge can be seen as distributing authorship: almost 30,000 buyers collectively determined the supply and composition of that work’s outcome, making the “most expensive NFT” (Courage Foundation, 2022) a product of crowd dynamics as much as Pak’s initial concept. This radical dispersal of agency recalls Umberto Eco’s idea of the “open work,” extended into the tokenized realm—the artwork is an evolving network of contributions. By building frameworks that require audience input to complete the work, Pak positions the community as co-creators—enacting what Zeilinger terms “structures of belonging” beyond mere ownership (Zeilinger, 2023). The artistic authorship becomes plural and emergent from the group, rather than originating solely from the individual artist.
@@ 4.4. Temporal dynamism and lifecycle aesthetics
Pak’s works are not static snapshots; they unfold over time, embracing change, ephemerality, and defined life cycles as core aesthetic elements. In classical systems art, the “consistency of a system may be altered in time and space” and all stages of its life cycle become material for the artwork. Pak’s projects exemplify this temporal dynamism. Lost Poets (2021) was structured as a narrative in multiple acts—from the initial sale of “Pages,” to a delayed Reveal where pages transformed into Poet NFTs, through daily drops of special “Origin” Poets, and onward towards a planned “Twist” and final Epilogue after 365 days. The piece lives and evolves across an entire year, its full meaning only emerging gradually as collectors engage over time. This built-in lifecycle—birth (minting), maturation (naming and interacting), and an end-state after one year (when burning for \$ASH is enabled)—turns time into an artistic medium.
@@
Other Pak works similarly rely on temporal structure. Censored’s Clock is explicitly a time-based artwork: a dynamic counter tethered to Assange’s real-world status—it tallied his days in prison until his release in June 2024, at which point it began counting days of freedom. The open edition portion of Censored was also time-limited to 48 hours, highlighting the urgency and performative moment of the audience’s participation. In X (2020), each NFT “moment” was only available during a one-day window, making “one single day of infinite existence” the defining limit of each edition. Scarcity and content in X were thus time-contingent: the fewer people who acted within the day, the fewer copies existed. Even Pak’s market dynamics often have temporal phases (for instance, The Fungible’s escalating price tiers each day). By designing pieces with evolving states and deadlines, Pak “recreates the cycle of life” in digital form, acknowledging that in a systemic artwork, impermanence and transformation are not byproducts but the very substance of the art.
4.5. Economy-driven engagement